Trust Company applies overhead based on direct labor hours. At the beginning of the year, Trust estimates overhead to be $700,000, machine hours to be 200,000, and direct labor hours to be 35,000. During February, Trust has 5,000 direct labor hours and 10,000 machine hours.

a.

What is the predetermined overhead rate?

b.

What is the amount of overhead applied for February?

c.

If the actual overhead for February is $98,300, what is the overhead variance, and is it overapplied or underapplied?
Answer #1

A) Predetermined overhead rate

= Estimated overhead/Estimated total labor hours

= 700,000/35,000

= 20 per dlh

B) Overhead applie

= 20 * 5,000

= 100,000

C) Over-applied overhead = 1,700

98,300 - 100,000 =